Journalist Report – February 26th

Hello Capcom !

Journalist report – Sol 08

Prepared by : Alexandre Martin (Crew Journalist) – 26 feb 2018

Here is my report of the day

Best regards.

Journalist report, 02/26/18: Water on Mars?

As most of the crew became increasingly hungry over the days, Benoît, with the help of Jérémy, decided to make pizzas. That was not an easy task using only dehydrated products! However, along with some seasonings that Jérémy brought back from the GreenHab, those pizzas were literally amazing, and to go to bed with a full stomach was deeply appreciated!

“Wake up”! This time, I was indeed not woken up by my alarm, that I forgot to put back after our day off, but by Victoria, knocking on my door. After a resting day for our muscles, it was time to go back to sport, and to bring back our body aches. An EVA was also planned for today. And for the first time of the mission, neither Louis nor Victoria would be EVA leader! Jérémy was the one to lead today, accompanied by Gabriel, Victoria and me. We made a short stop at our experiments, in order to change the LOAC battery. However, the voltmeter seemed to have caught a cold, and could not give any information. Jérémy thus almost confused the battery to take back to the Hab with the battery to plug to the LOAC. After having gone further and further north since the beginning of the mission, we decided to travel south this time. However, our map played a few tricks to us, and we had some troubles finding the place we wanted to go to.

We finally arrived at the rim of the canyon we wanted to visit, and we made a small detour in order to be able to get into it. We discovered here something we did not think we would find in such environments: some large bodies of frozen shallow water. Jérémy could not prevent himself from walking on one of them, yet he rapidly stopped when he realized that some cracks began to initiate. After having completed our trip in the canyon and having taken some photos, we went back to our vehicles, and we decided to go back home. This has been a quite short yet very nice EVA!

As our reserves of pasta and rice are starting to decline dangerously, we now have to assault our stock of beans, surely nourishing, but still a lot less tasty. As usual, we spent a lot of our afternoon time playing “Keep talking and nobody explodes”: the difficulty is now way higher (Louis quite notably paid the price) and our team have been split up, leading to some misapprehension and misunderstandings.

Alexandre Martin, Crew 189 Journalist

Operations Report – February 26th

Hello CapCom, Graeme,

Here is my Operation Report.

Crew 189 Operations Report 26Feb2018

SOL: 8

Name of person filing report: Gabriel PAYEN

Non-nominal systems: Pipe leak, radio

Notes on non-nominal systems:

-The pipe still leaks in the lower deck

-A radio became inaudible during the EVA

Generator (hours run): Turned off this morning and turned on at 6:45 PM

Solar— SOC % (Before generator is run at night) 85%

Diesel – Too dark to see

Propane – Too dark to see

Ethanol Free Gasoline (5 Gallon containers for ATV) – 5 gallons.

Water (trailer) – 0 gallons.

Water (static) – Between 400 and 450 gallons

Trailer to Static Pump used – NO

Water (loft) – Static to Loft Pump used – YES

Water Meter: 131 473.9

Toilet tank emptied: NO

ATV’s Used: 350.1 & 350.2

Oil Added : YES

ATV Fuel Used: Gals

Hours the ATVs were Used today: 3 hours

Notes on ATVs: –

Deimos rover used: NO

Hours:

Beginning charge:

Ending charge:

Currently charging:

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours:

Beginning charge:

Ending charge:

Currently charging:

Spirit rover used: NO

Hours:

Beginning charge:

Ending charge:

Currently charging:

Opportunity rover used: NO

Hours:

Beginning charge:

Ending charge:

Currently charging:

Curiosity rover used: NO

Hours: Beginning of the EVA : 13.3 , End of the EVA : 13.7

Beginning charge: 100%

Ending charge: 87%

Currently charging: YES

HabCar used and why, where : –

General notes and comments: –

Summary of internet: –

Summary of suits and radios:

For the first time, our helmets stayed clear during the EVA and the fog quickly disappeared after every physical effort!

Summary of Hab operations: –

Summary of GreenHab operations: –

Summary of ScienceDome operations: –

Summary of RAMM operations: –

Summary of health and safety issues: –

Questions, concerns and requests to Mission Support: –

Sol Summary – February 26th

Hello again Capcom,

Here is today’s sol summary:

Crew 189 Sol Summary Report 26 Feb 2018

Sol 8

Summary Title: Getting the rhythm

Author: Louis Mangin

Mission Status: All nominal

Sol Activity Summary:
Today started by a quick EVA, to change LOAC’s battery, followed by an exploration of White Canyon. We spent this afternoon working on experiments and cleaning up the Hab as lower deck was becoming very dirty. The crew is now in good shape and the three of us recording our physical performance this morning scored new records in almost every discipline.
Only some of us lack of sleep, but except from this, everybody is doing great! The only thing worrying us is reduced food’s portions, but we start to adapt to lighter meals.

Look Ahead Plan: Emergency procedures training in the morning, in situ testing during a quick EVA near the hab in the afternoon.

Anomalies in work: Water leak in the Hab Lower deck

Weather: Mostly sunny.

Crew Physical Status: Fine.

EVA: –

Reports to be file:
Commander
Journalist
Engineering
Greenhab Officer
EVA report
HSO report

Support Requested: –

Louis Mangin
Crew 189 Commander

Commander Report – February 26th

Hi CapCom !

Here is the Crew 189 Commander Report 26Feb2018

Sol 8

Title: Beginning of our second week on Mars

Dear Earth,

We are currently beginning our second week on Mars. Our Sunday day-off was needed and everyone seemed to enjoy it. As a celebration, Benoit cooked the ideal meal to end at best our first week on Mars: homemade pizzas from our dry food (homemade dough, tomato sauce, cheese, meat, and fresh herbs from our greenhab).

From my commander view, this simulation is a success for now: It is like it started really good and just kept getting better. After the debriefing of our physical training results that Louis gave us yesterday, we are all ready and motivated to meet the highest expectations of our sports coach. Benoit, Laurent and Louis did their best scores (let’s see tomorrow if the other part of the team continues to be as successful).

After our breakfast, I let Jérémy organizing the EVA preparation as he was EVA leader. Because of some backpacks issues, we were 10 minutes late. The EVA team composed of Jérémy, Gabriel, Alex and I went out to the Pooh’s corner place where our two air sensors were deployed last week. Jérémy changed the battery of his experiment. After that, we took two ATVs and a rover to explore and take pictures at White Rock Canyon, which I discovered last year and remembered well. Meanwhile, Laurent explained Louis the emergency procedures he wants to simulate during future EVAs. The EVA team came back earlier and we managed to check and clean all our EVA equipment before lunchtime. We cooked beans with meat and enchiladas sauce (we have beans in large quantity, beside all the junk snacks).

The second week begins with a big concern: the crew is more and more hungry. During the first week, the protein bars have been quickly eaten and we decided today to give everyone the same numbers of the remaining bars: who is going to be the first to eat all of them? (I bet on Benoit or Jérémy)

After a quick break, we “played” again the Human Factors experiment of Gabriel. This week the competition between the two teams is increasing because the better team will win a surprise box with food after the simulation! Everyone seems very excited to take part in this experiment, and we can’t wait for the competition results (and even more because with my friend Louis we will lead the two different teams).

After the delicious meal of yesterday, we are beginning to talk about our food rationing for the next two weeks. The success of our bread competition made the flour stocks cruelly decreasing… I let you guess what will be our future meals (it starts with B and ends with EANS)

Ad Astra!

Victoria Da-Poian

Commander of the beans-eaters Crew 189

EVA Report – February 26th

Dear CapCom,

Here is the EVA Report:

26 Feb 2018

EVA #6 Crew members: Jérémy Auclair (EVA leader), Victoria Da-Poïan, Gabriel Payen, Alexandre Martin

Location:

– Near Pooh’s Corner: 12S-518900E-4252000N

– White Rock Canyon: 12S-520200E-4247500N

Time: departure at 09:09 a.m. Duration: 2 hour 13 minutes

Narrative: My first EVA as EVA leader went really well ! We changer the battery that powered the LOAC instrument (the old one was completely empty), we tested the battery powering the MegaARES instrument and confirmed it needed to be changed quickly (tomorrow). Otherwise both instruments are supporting the desert conditions perfectly. We then went to white rock canyon for exploration and photography. The walk went perfectly, the temperature was perfect, our reparations on the back packs prevented the apparition of fog on our helmets. We started to walk inside the canyon and on our way back we walk on the ridge (at a safe distance) to change the point of view. We finished the walk around 11am. Having nothing else to do we went back to the Hab to refill the ATVs and end the EVA. Everything went fine.

Purposes:

– LOAC: Maintenance – replacement of the battery

– MegaARES: Maintenance – Battery level test

– White Rock Canyon – Exploration and photography

Vehicles: 1 Rover (Curiosity), 2 ATVs (350.1, 350.2)

9:09 a.m.: EVA#6 left Airlock

9:15 a.m.: EVA left the Hab using Curiosity, ATVs 350.1 and 350.2

9:23 a.m.: Arrival on the working place

9:25 a.m.: Beginning of LOAC battery replacement and MegaARES battery test

9:35 a.m.: End of maintenance

9:40 a.m.: Beginning of the drive for White Rock Canyon

10:00 a.m.: Arrival at White Rock Canyon and beginning of walk

10:50 a.m.: End of walk and beginning of the drive back to the Hab

11:05 a.m.: Arrival at the Hab

11:15 a.m.: Entry in the Airlock

Best,

Victoria

Commander of Crew 189

Greenhab Report – February 26th

Dear CapCom and Graeme,

Here is the GreenHab Report

AUCLAIR Jérémy

26 feb 2018

Environmental control:

(door open around 11am, left closed after 14h)

Heating once the sun was down

80 % Shade cloth on

Average temperatures:

Around 35 °C at 11am, around 28 °C the rest of the day

Hours of supplemental light: from 7pm to 9pm

Changes to crops: Lettuce baby plants planted in bigger pots

Daily water usage for crops: around 12 gallons

Time(s) of watering for crops: 11:30 am and 7:30pm

Narrative: Still achieving the same type work

Support/supplies needed: wooden sticks to support falling plants

Cheers,

Victoria

Commander of Crew189